OpenStudy (anonymous):

if you went to the store to buy a sweater marked $70, and it was on 40% off rack, how much is the sweater before tax?

OpenStudy (anonymous):

.6*$70

OpenStudy (anonymous):

100% = x $ 60% = 70 $ //* 1,66 100% = x$ 100% = 70* 1,66 = 70 * 5/3 = 350 /3 ~ 116,67

OpenStudy (anonymous):

how much after a 7% sales tax is added on?

OpenStudy (anonymous):

after the 40% off ?

OpenStudy (anonymous):

70$ * 1.06 = 74.2$